I am not understanding sequences and series? Can someone help me understand this better and show the steps. Solve question given that a3 = 3 then a6= -81 Find r? Find a1. Thank you

Answer by stanbon(75887)

given that a3 = 3 then a6= -81 Find r? Find a1
-----------------------------
Since you are looking for "r" your series is a geometric series.
-------------------------
a3 = a1*r^2 = 3
a6 = a1*r^5 = -81
--------------------------
Divide a6 by a3 to get:
r^3 = -27
r = -3
----------------
Substitute r = -3 into a3 to solve for a1
a1*(-3)^2 = 3
a1 = 3/9 = 1/3
-------------------------
Cheers,
Stan H.